Supply = -20+10p
Demand = 400-20p
Solve for equilibrium price and quantity
1
Expert's answer
2019-10-22T09:16:18-0400

To calculate equilibrium price equate the supply equation to the demand equation and solve for p


"-20+10p=400-40p"

"50p=420"


"p=8.40"

To get equilibrium quantity substitute 8.40 for p in either equation


"400-40\times8.4=64"